OpenStudy (bookworm14):

Please check my answer !!! The base of a square pyramid has a side length of 7 inches. The height of the pyramid is 19 inches. What is the approximate volume of the pyramid? A. 133 in3 B. 177 in3 C. 310 in3 D. 931 in3 I think it is D. but i'm not sure

OpenStudy (anonymous):

Hmm..

OpenStudy (anonymous):

I Got D.. But I'm Not The Smartest Person Out There :p

OpenStudy (anonymous):

I pretty sure its A

OpenStudy (bookworm14):

thats what i thought, however they told me to use the formulas V=s^3 or 1/3 Bh not really sure ??

OpenStudy (anonymous):

Hold Up.. I'mma Check

OpenStudy (bookworm14):

i got D somehow

OpenStudy (bookworm14):

can you please explain how, ?

OpenStudy (anonymous):

Yea I Got A..

OpenStudy (anonymous):

Accoroding to the\[2 \pi \times r \times h= 2\pi \times3.5\times19=133\]

OpenStudy (bookworm14):

can you use the formulas i gave you and get the same answer ?

OpenStudy (anonymous):

what formula?

OpenStudy (bookworm14):

V=S^3 , or 1/3Bh

OpenStudy (bookworm14):

nothin ?

OpenStudy (anonymous):

Do you got it

OpenStudy (math_genius123):

It should be 310 because its approximately 310.33 then just round you get 310

OpenStudy (math_genius123):

so C.

OpenStudy (math_genius123):

First find the base area\[B=s ^{2}=7^{2}\] Then the equation to find volume \[V=\frac{ 1 }{ 3 }Bh\]\[=\frac{ 1 }{ 3 }\times49\times19\]\[=310.333333333\]Then you round so it will be\[310\]